OpenStudy (anonymous):

What is the line of symmetry for the parabola whose equation is y = x^2 + 10x + 25?

OpenStudy (anonymous):

x= -10?

OpenStudy (anonymous):

you can write this as \[\Large y=(x+5)^2=(x-(-5)^2\] so line of symmetry should be x=-5